OpenStudy (anonymous):

how can u simplify 1-x3/ 1-x

OpenStudy (anonymous):

= x^2+x+1

OpenStudy (anonymous):

ix this ( 1 - x^3) / (1 - x) ?

OpenStudy (anonymous):

factor thenumerator using two cubed differnce identity then cancel with the denomerator

OpenStudy (anonymous):

yes it is 1-x^3/ 1-x

OpenStudy (anonymous):

then raheen's answer is correct

OpenStudy (anonymous):

how did he solve it

OpenStudy (anonymous):

well 1 - x^3 = (1-x)(x^2 + x +1) and the 1-x cancels with the 1-x on the top leaving only the second bracket x^2 + x +1)

OpenStudy (anonymous):

thank u :)

OpenStudy (anonymous):

yw